# Learning Mixtures of Linear Dynamical Systems (MoLDS) via Hybrid Tensor–EM Method

### v1

@@ D.2 PMd Dataset
The dorsal premotor cortex (PMd) dataset contains single–trial reaches, where time–aligned kinematics (position/velocity/acceleration) and multi–unit spikes, plus a reach direction angle $\theta\in(-\pi,\pi]$ are stored. For MoLDS, the input and observation are taken as
@@
where $\mathrm{speed}=\sqrt{v_{x}^{2}+v_{y}^{2}}$ and $Y_{t}$ denotes PCA–reduced, zz–scored firing rates. As in Area2 (App. D.1), angles are used only for analysis/visualization; MoLDS is trained without angle labels.
@@
Analysis.
We reuse the analysis logic from Area2. As the direction angles are not discrete as in Area2, we split them into 12 bins, and each bin is colored by the dominant component (largest mean $r_{ik}$ in the bin). Component–wise preferred directions are shown by responsibility–weighted arrows
@@
and plot (a) the compact magnitude curves $\|g_{k}(\tau)\|_{F}$ and (b) per–input energies $E_{k,j}(\tau)=\|g_{k}(\tau)_{:,j}\|_{2}$ for $j\in\{v_{x},v_{y},a_{x},a_{y},\mathrm{speed}\}$. For a global view, we also vectorize $\{g_{k}(\tau)\}_{k,\tau}$, run PCA, and scatter the first two PCs.

### v2

@@ E.2 PMd Dataset
The dorsal premotor cortex (PMd) dataset contains single-trial reaches, where time-aligned kinematics (position/velocity/acceleration) and multi-unit spikes, plus a reach direction angle $\theta\in(-\pi,\pi]$ are stored. For MoLDS, the input and observation are taken as
@@
where $\mathrm{speed}=\sqrt{v_{x}^{2}+v_{y}^{2}}$ and $Y_{t}$ denotes PCA-reduced, $z$-scored firing rates. Here we use $p=16$ PCs, which explain $90\%$ of neural activity. As in Area2 (App. E.1), angles are used only for analysis/visualization; MoLDS is trained without angle labels.
@@
Analysis idea.
We reuse the analysis logic from Area2. As the direction angles are not discrete as in Area2, we split them into 12 bins, and each bin is colored by the dominant component (largest mean $r_{ik}$ in the bin). Component-wise preferred directions are shown by responsibility-weighted arrows
@@
and plot (a) the compact magnitude curves $\|g_{k}(\tau)\|_{F}$ and (b) per-input energies $E_{k,j}(\tau)=\|g_{k}(\tau)_{:,j}\|_{2}$ for $j\in\{v_{x},v_{y},a_{x},a_{y},\mathrm{speed}\}$. For a global view, we also vectorize $\{g_{k}(\tau)\}_{k,\tau}$, run PCA, and scatter the first two PCs.